# DDSC: Dynamic Dual-Signal Curriculum for Data-Efficient Acoustic Scene Classification under Domain Shift

### v1

@@ 1 Introduction
To mitigate domain shift with limited annotations, prior work [4, 6, 7, 8] explores Curriculum Learning (CL) to encourage domain-invariant representations. CL mimics the human progression from easy to hard and structures the presentation of training data [9], facilitating the acquisition of domain-invariant, generalizable representations [10].Following this paradigm, these methods predefine an easy-to-hard curriculum that reorders or reweights training examples, encouraging the model to first focus on domain-invariant samples and later incorporate device-specific ones.
@@
To this end, we propose the Dynamic Dual-Signal Curriculum (DDSC)  [^1] Code will be released upon acceptance., a dynamic curriculum that adaptively adjusts sample weights. At each epoch, DDSC computes two complementary signals: (i) a domain-invariance score obtained from prototype–posterior entropy over device prototypes, and (ii) a learning-progress score derived from the smoothed change of the per-example loss across epochs. A time-varying scheduler fuses these signals into a continuous curriculum score that determines per-example weights. The resulting schedule realizes an easy-to-hard presentation—prioritizing domain-invariant examples in early epochs and progressively emphasizing device-specific, difficult cases—thereby improving generalization to unseen devices without architectural changes or inference overhead.
@@ 2.4 Signal Fusion and Weighted Objective
where $\lambda_{\min}=0.2$ was selected through validation on the DCASE2024 offical baseline. This value was found to provide stable performance across all tested models and label budgets.

### v2

@@ 1 Introduction
To mitigate domain shift with limited annotations, prior work [20, 18, 16, 21] explores Curriculum Learning (CL) to encourage domain-invariant representations. CL mimics the human progression from easy to hard and structures the presentation of training data [3], facilitating the acquisition of domain-invariant, generalizable representations [9, 19].Following this paradigm, these methods predefine an easy-to-hard curriculum that reorders or reweights training examples, encouraging the model to first focus on domain-invariant samples and later incorporate device-specific ones.
@@
To this end, we propose the Dynamic Dual-Signal Curriculum (DDSC), a dynamic curriculum that adaptively adjusts sample weights. At each epoch, DDSC computes two complementary signals: (i) a domain-invariance score obtained from prototype–posterior entropy over device prototypes, and (ii) a learning-progress score derived from the smoothed change of the per-example loss across epochs. A time-varying scheduler fuses these signals into a continuous curriculum score that determines per-example weights. The resulting schedule realizes an easy-to-hard presentation—prioritizing domain-invariant examples in early epochs and progressively emphasizing device-specific, difficult cases—thereby improving generalization to unseen devices without architectural changes or inference overhead.
@@ 2.4 Signal Fusion and Weighted Objective
where $\lambda_{\min}=0.2$ was selected through validation on the DCASE2024 official baseline. This value was found to provide stable performance across all tested models and label budgets.
@@ 4 Conclusion
5 Acknowledgment
This work was funded by Basic Research Program of Jiangsu Province under Grant BG2024027.